OpenStudy (anonymous):

logx-log4=log12 i have no idea how to do this! please help!

OpenStudy (anonymous):

Use formula loga - logb = log(a / b) So we have log(x/4)= log(12) now cancel log from both side s we get x/4 = 12 so just solve for x

OpenStudy (anonymous):

oh wow thanks you're amazing!!

OpenStudy (anonymous):

u r welcome :)